OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 12x = -5

OpenStudy (anonymous):

Completing the square step by step step 1 : use b = 12 where b = coefficient of x in this case b=12 Step 2 : add (b/2)^2 to both sides in this case (b/2) = (12/2)^2 = 6^2 =36 step 3 : the left side is a perfect square So : X^2 +12x= -5 X^2+12X+ 36 = -5+36 = 31 (x+6)^2 =31 Answer is (b/2)^2 =36